OpenStudy (anonymous):

factor this -y^2+17y-72

OpenStudy (anonymous):

\[(-y ^{2}+17y)-72 = -(y ^{2}-17y)-72 = -(y ^{2}-17y+289/4)-72+289/4\] \[= -(y-17/2)^{2}+1/4\]